# Generative AI-Powered Plugin for Robust

### v1

@@ IV-B1 Impact on AI Model Diversity
Figure 5: Performance comparison of text classification models with and without our proposed plugin. Results demonstrate significant accuracy improvements across all models, with particularly notable gains in domain-specialized models (BioBERT, BioMedBERT) compared to their baseline performance.
In this experiment, we examined how our proposed plugin enhances the performance of various text classification models in FL environments. We compared performance improvements for four different models: general-purpose models BERT [22] and DistilBERT [23], and medical domain pre-trained models BioBERT [24] and BioMedBERT [25]. Through this experimental setup, we aimed to demonstrate our plugin’s capability to enhance model performance regardless of their initial domain specialization. Additionally, for data augmentation in Phase 1, multiple LLMs were utilized to generate synthetic data, ensuring balanced representation across all edge devices.
Performance evaluation revealed significant improvements across all models when applying our proposed plugin. As shown in Fig. 5, BERT’s accuracy increased from 29.8% to 47.3% and DistilBERT improved from 32.2% to 46.6%. Most notably, medical domain pre-trained models showed even more substantial gains, with BioBERT’s accuracy rising from 42.1% to 53.4% and BioMedBERT reaching 52.0% from a baseline of 28.7%. These significant improvements demonstrate our plugin’s effectiveness in enhancing model performance while preserving the advantages of domain-specific pre-training. The results particularly highlight how our plugin can amplify the benefits of domain-specialized models in federated medical applications. Based on these results, we use BioBERT as the edge device AI model for subsequent experiments.
Detailed analysis revealed the synergistic effect between our plugin and domain-specialized models. The performance gains were particularly pronounced for medical domain pre-trained models, with BioBERT showing 7.7% and 6.8% higher accuracy compared to BERT and DistilBERT respectively when enhanced by our plugin. These improvements demonstrate that our synthetic data augmentation and balanced sampling strategies effectively complement domain-specific knowledge, creating more robust FL systems. Furthermore, the significant performance enhancements across all models validate our plugin’s versatility in addressing data heterogeneity challenges.
@@ IV-B2 Impact on Various FL Methods
Figure 6: Performance comparison of different FL algorithms enhanced with the proposed plugin. Results demonstrate consistent improvements in test accuracy across FedAvg, FedProx, and FedRS, with various LLM configurations, highlighting the plugin’s effectiveness and versatility.
To demonstrate the versatility and effectiveness of our proposed plugin across different federated optimization techniques, we designed comparative experiments. Our evaluation framework considered three key aspects: the plugin’s adaptability to various FL algorithms, its impact on convergence speed, and overall accuracy improvements. Through this comprehensive evaluation setup, we aimed to validate that our plugin can enhance performance across different FL approaches.
Performance analysis revealed consistent improvements across all FL algorithms when combined with our proposed plugin. Among tested combinations, FedRS with Llama3.0:8B achieved the highest accuracy at 53.67%, while the same configuration with FedProx and FedAvg reached 53.51% and 53.42% respectively as shown in Fig. 6. Notably, all three FL algorithms showed significant performance gains compared to their baseline implementations, with improvements ranging from 9% to 11%. This consistent enhancement across different FL approaches demonstrates the plugin’s effectiveness in mitigating data heterogeneity through synthetic data augmentation and balanced sampling.
Convergence analysis showed that all configurations with our plugin reached stable performance significantly faster than baseline approaches. While baseline FL algorithms required 23 epochs to reach maximum accuracy, configurations with our plugin consistently achieved convergence between epochs 15-20. Specifically, Gemma 2.0:9B, Llama 3.2 series, and Phi 2.0:3B demonstrated rapid convergence at epoch 15 across all FL algorithms, showing approximately 35% reduction in required training epochs. These results suggest that our proposed plugin effectively accelerates convergence regardless of the underlying FL algorithm or LLM choice. The experimental outcomes demonstrate that our synthetic data augmentation and balanced sampling strategies not only improve accuracy but also significantly enhance training efficiency across different FL frameworks.
IV-B3 Impact of Data Heterogeneity
Figure 7: Performance comparison across different degrees of data heterogeneity $\alpha$. Results demonstrate the plugin’s effectiveness in maintaining high accuracy even under extreme Non-IID conditions, with consistent improvements observed across various LLM configurations.
To validate our plugin’s effectiveness under different levels of data heterogeneity, we conducted experiments with varying degrees of Non-IID distribution controlled by the concentration parameter $\alpha$. Lower values of $\alpha$ indicate higher data heterogeneity among edge devices, presenting more challenging FL scenarios. This experimental setup particularly focuses on extreme Non-IID situations where conventional federated learning approaches often struggle to maintain performance.
Performance analysis revealed that our proposed plugin demonstrates increasingly significant improvements as data heterogeneity increases. With Gemma 1.0:7B, the plugin maintained 38.89% accuracy even at $\alpha=0.01$, while baseline FedAvg accuracy dropped to 21.43% as shown in Fig. 7. Similar patterns were observed across different LLM families, with Gemma 2.0:2b and Gemma 2.0:9b achieving 48.13% and 47.68% accuracy respectively at high heterogeneity levels. These results indicate that synthetic data generation effectively compensates for data imbalances, particularly in scenarios with minimal data overlap between devices.
Figure 8: UMAP visualization of embedding vectors from original and synthetic data across different LLM configurations. Distinct clustering patterns reveal varying capabilities of LLMs in preserving the original data distribution, with Phi 3.0 and 3.5 showing notable deviations from the source distribution.
Notably, under extreme Non-IID conditions $\alpha\leq 0.001$, our plugin consistently provided substantial performance improvements compared to unused case. The enhanced performance remained stable across all tested LLM configurations, with accuracy improvements ranging from 11% to 28% at $\alpha=0.001$. These findings demonstrate that our plugin effectively addresses one of the fundamental challenges in FL: maintaining model performance under severe data heterogeneity. The robust performance across different heterogeneity levels suggests strong potential for real-world applications where Non-IID data distribution is inevitable.
IV-B4 Analysis of Synthetic Data Distribution
To understand the effectiveness of different LLMs in generating representative synthetic data, we conducted an embedding space analysis. We visualized the distribution of both original and synthetic samples using UMAP [26] dimensionality reduction technique as shown in Fig. 8. Embedding vectors were extracted using ’MXBai-Embed-Large model’, providing a two-dimensional representation of the high-dimensional text features. This visualization approach enables direct comparison of the semantic relationships between original medical texts and their synthetic counterparts. Such comparisons offer insights into each LLM’s capability to maintain the underlying data distribution.
Comparative analysis revealed distinct patterns across different LLM families in preserving the original data distribution. Among them, Gemma 2.0:9B and Llama 3.2:3B demonstrated particularly strong alignment between synthetic and original data clusters, maintaining clear separation between different medical text categories while preserving intra-class relationships. By successfully capturing the inherent structure of the medical text data, these models validated their superior performance observed in previous experiments. Additionally, the synthetic samples generated by these models showed consistent overlap with original data clusters, providing clear evidence for their effective preservation of domain-specific features.
Notably, Phi 3.0:3B and Phi 3.5:3B exhibited significant deviation from the original data distribution. Their synthetic samples formed distinct clusters that were clearly separated from the original data points, unlike other LLM families. This distribution mismatch explains their relatively lower performance in previous experiments, particularly in maintaining classification accuracy. This misalignment between original and synthetic data distributions particularly impacts the effectiveness of our plugin’s data augmentation strategy. These findings ultimately highlight the importance of selecting appropriate generative models that can maintain semantic consistency for synthetic data generation.
@@ V Conclusion
In this paper, we proposed a novel plugin for FL that addresses data heterogeneity challenges through generative AI-enhanced data augmentation and balanced sampling strategy. Our experiments demonstrated the plugin’s effectiveness across various text classification models and FL algorithms while reducing required training epochs. The plugin showed strong performance improvements in medical domain-specialized models, with BioBERT achieving 53.4% accuracy from a 42.1% baseline. Our approach maintained robust performance under extreme Non-IID conditions, showing consistent improvements across different LLM configurations. These results establish our plugin as a flexible solution for enhancing FL performance in real-world heterogeneous environments.

### v3

@@ IV-B1 Impact on AI Model Diversity
Figure 4: Performance comparison of text classification models with and without our proposed plugin.
In this experiment, we evaluated our plugin’s ability to enhance text classification models in FL environments. Moreover, we compared four models: general-purpose BERT [kenton2019bert] and DistilBERT [sanh2019distilbert], and domain-specific BioBERT [lee2020biobert] and BioMedBERT [chakraborty2020biomedbert]. As shown in Fig. 4, performance evaluation showed significant improvements, with BERT’s accuracy increasing from 29.8% to 47.3% and DistilBERT’s from 32.2% to 46.6%. In addition, the medical domain pre-trained models achieved even greater gains, with BioBERT rising from 42.1% to 53.4% and BioMedBERT from 28.7% to 52.0%. Thus, based on these results, we use BioBERT as the edge device model for the subsequent experiment.
@@ IV-B2 Impact on Various FL Methods
Figure 5: Performance comparison of different FL algorithms enhanced with the proposed plugin.
To demonstrate the versatility of our proposed plugin, we examined its impact on convergence speed and overall performance across various FL optimization methods. As shown in Fig. 5, the FL methods enhanced with our plugin reached stable performance significantly faster than their vanilla counterparts. In fact, while the vanilla FL methods required 23 epochs to attain maximum accuracy, configurations with our plugin consistently converged between epochs 15 and 20. Notably, models such as Gemma 2.0:9B, the Llama 3.2 series, and Phi 2.0:3B achieved convergence as early as epoch 15, reflecting an approximate 35% reduction in training epochs. Therefore, the results indicate that the proposed synthetic data augmentation and balanced sampling strategy improves both the convergence efficiency and global accuracy in FL.
@@ V Conclusion
In this paper, we proposed a novel plugin for FL methods that addresses data heterogeneity challenges through generative AI-enhanced data augmentation and balanced sampling strategy. In particular, the plugin balances local data distributions through generative augmentation and selects statistically representative devices for aggregation, improving robustness. The plugin showed strong performance improvements in medical domain-specialized models, with BioBERT achieving 53.4% accuracy from a 42.1% baseline. Moreover, we numerically demonstrated the plugin’s effectiveness across FL methods while reducing required training epochs. These results establish our plugin as a flexible solution for enhancing FL performance in real-world heterogeneous environments. As future work, we plan to extend this plugin to multimodal FL settings to further broaden its applicability.